# Multi-modal Retrieval Augmented Multi-modal Generation: A Benchmark, Evaluate Metrics and Strong Baselines

### v1

@@ 4.3 Element Evaluation
As described in Section 3, we need to perform In-Doc retrieval to select text and visual elements for efficient generation. To achieve this goal, the relevance between each element and the user query should be evaluated. Specifically, we utilized the DeepSeek-Chat DeepSeek-AI (2024) API for evaluating textual elements and MiniCPM-V-2.6 Yao et al. (2024) for assessing visual elements. Our preliminary manual annotations revealed that these methods demonstrate a high degree of consistency with human evaluations in measuring correlation, making them suitable for this task [^5] Please refer to Appendix B for more details.. This step results in a score from 0 to 10 for each element, where higher scores indicate better relevancy. This can be used to replace the traditional embedding-based similarity score for retrieval.
With the above procedures, a high-quality multi-modal RAG dataset with real-world user queries, retrieved web pages and images is constructed. Detailed statistical information of our constructed dataset can be found in Appendix C.
5 Strong Baselines for M2RAG Task
Finishing Step 1-3 of Figure 2, we introduce three strong baselines for Step 4 to accomplish our proposed M2RAG task.
5.1 Separate Modeling
Previous studies have approached text generation and image integration as distinct processes, exemplified by the MuRAR method Zhu et al. (2024). MuRAR operates in three sequential steps: (1) retrieving relevant plain text content and generating a text-based response; (2) identifying and incorporating appropriate visual elements into the text-based response; and (3) refining the text to seamlessly integrate the visual elements and their associated information.
In our implementation, pre-trained embedding model Reimers and Gurevych (2019) is used to retrieve visual elements based on their surrounding text content or detailed image caption generated by GPT-4o. Since the separated method does not need to process multi-modal content, it is implemented by LLMs.
5.2 Joint Modeling
The separate approach overlooks the relationships between textual content and multi-modal data when generating text responses, leading to sub-optimal integration of multi-modal information. To address this limitation, we propose several baseline methods that jointly model text and images as a unified entity, capturing their rich relationships and interactive dynamics.
Specifically, our approach can be categorized into single-stage and multi-stage methods, leveraging LLMs or MLLMs as backbone models. This design allows for a more holistic representation of multi-modal content, as depicted in Step 4 of Figure 2.
Single- and Multi-stage Approaches
@@ 5.2 Joint Modeling
Single-stage approach requires models to directly generated a multi-modal output with selected images in their corresponding placeholders, given all multi-modal content saved in a single prompt.
@@
Although the single-stage approach is straightforward, current open-source foundation models struggle in generating accurate image placeholders. Besides, it is a well-known limitation that most MLLMs may struggle with handling a large number of images simultaneously. Therefore, we propose a multi-stage approach to address the complexity of the single-stage approach, which involves three stages: (1) Text Response Generation: generate a plain text answer given the multi-modal input; (2) Image Interleaving: divide the text into several segments, and LLMs or MLLMs are prompted to determine which segment requires the inserted images for better readability; (3) Text Refinement: refine each segment by incorporating the information of selected images. Although multi-stage approach addresses the shortcomings of single-stage approach, their computational overhead for inference is much higher.
LLM and MLLM Backbones
Both of LLMs and MLLMs can be used to solve M2RAG task with single-stage or multi-stage approaches. Specifically, as introduced in Section 3, given the retrieved multi-modal data resource $R_{\text{In-Doc}}$, the prompt $P$ for LLMs and MLLMs could be constructed using a structured template $T$:
$P=T(G,Q,R_{\text{In-Doc}})$ (2)
where $G,Q$ denotes the textual task guidelines and user query, respectively. Then LLMs and MLLMs could be prompted with $P$ to solve M2RAG task: (1) For MLLMs, the encoded image tokens are placed at their original positions  Dubey et al. (2024); Wang et al. (2024a), reserving the relative position in the multi-modal web pages. Since the additional images do not belong to any web pages, their image tokens are placed in a distinct section in the prompt; (2) For LLMs, since they could not process images directly, the detailed captions of images in $R_{\text{In-Doc}}$ are provided, allowing LLMs to access the semantical content of images.
6 Evaluation Metrics
To evaluate the effectiveness of numerous models and methods, we design four text-modal metrics and four multi-modal metrics. Since the generated response are open-ended, these evaluation metrics are evaluated by prompting advanced GPT-4o model in a reference-free manner Lan et al. (2024). To the best of our knowledge, prompting GPT-4o for automatic evaluation is the most reliable automatic evaluation methods for open-ended RAG and multi-modal evaluation scenarios Es et al. (2023); Ku et al. (2024); Lu et al. (2023). The detailed prompts for our proposed text-only and multi-modal evaluation metrics can be found in Appendix D. The score ranges of all fine-grained text-only and multi-modal metrics are from 0 to 1, and scores close to 1 indicates better performance.
6.1 Text-modal Metrics
Following previous works Zhu et al. (2024), we utilize both linguistic and RAG metrics Es et al. (2023) to assess the quality of the generated content in the M2RAG task. These metrics ensure a comprehensive evaluation of the generated textual content.
Fluency
assesses the linguistic quality of model-generated text, ensuring the grammatical correctness, coherence, and readability Su et al. (2022); Es et al. (2023).
Relevance
assesses how pertinent the model-generated textual content is to the given user query Lan et al. (2022); Es et al. (2023).
Context Precision
evaluates the proportion of relevant chunks in the retrieved web pages by analyzing the overlap of their key concepts Es et al. (2023).
Faithfulness
measures the accuracy of the model-generated content in representing information from the original materials, focusing on factual alignment to avoid fabricated or misleading details Es et al. (2023); Sun et al. (2024).
6.2 Multi-modal Metrics
In addition to text-only metrics, we also focus on the quality of model-generated multi-modal content from following five perspectives.
Image Coherence
examines the logical and coherent alignment of images with their surrounding text content, ensuring visual content enhances and complements the textual narrative Zhu et al. (2024).
Image Helpfulness
evaluates how effectively images contribute to the user’s comprehension of the text. This involves assessing whether visuals provide additional insights, clarify complex ideas, or support textual details Zhu et al. (2024).
Image Reference
verifies the appropriateness and alignment between images and their textual references. Inadequate or incorrect references of images lead to the confusion during reading.
Image Recall
measures whether the generated content includes highly relevant, informative, and important images. As described in Section 4.3, images are evaluated and scored during pre-processing, with selection based on these scores prior to generation. Consequently, the images input to generative models are generally relevant to the user query. However, issues such as semantic overlap or quality variations among images may arise, leading to the inclusion of sub-optimal images in the retrieved images. The goal of Image Recall is to ensure that no images other than the sub-optimal ones are excluded from the final answer. To achieve this, we prompt MLLMs to assess whether each image enhances human understanding of the content. To address potential semantic overlap, we include textual descriptions of images already present in the document and instruct the model to classify an image as negative if its content is significantly redundant compared to existing images. This process ensures that only diverse and meaningful images are retained.
6.3 Overall Score Computation
Except for these fine-grained evaluation dimensions, it is essential to summarize a final overall score reflecting the capabilities of foundation models in M2RAG task. Specifically, the average of four text-only metrics and five multi-modal metrics is computed as the overall score.
@@ 7 Experimental Results
Model Type Approach Model Text-modal Metrics Multi-modal Metrics Overall
@@
LLMs Separate GPT-4o 0.80 0.82 0.69 0.89 0.53 0.44 0.43 0.43 0.63
Single GPT-4o 0.79 0.81 0.76 0.91 0.69 0.57 0.31 0.57 0.68
Llama-3.1-70B-Instruct 0.75 0.79 0.75 0.85 0.57 0.47 0.18 0.22 0.57
Qwen2.5-72B-Instruct 0.79 0.83 0.77 0.87 0.62 0.52 0.35 0.52 0.66
Llama-3.1-8B-Instruct 0.74 0.77 0.72 0.83 0.49 0.38 0.28 0.62 0.60
Qwen2.5-7B-Instruct 0.71 0.73 0.73 0.91 0.52 0.41 0.22 0.63 0.61
Average 0.75 0.79 0.75 0.87 0.56 0.45 0.27 0.51 0.62
Multi GPT-4o 0.78 0.76 0.68 0.87 0.68 0.60 0.81 0.97 0.77
Llama-3.1-70B-Instruct 0.73 0.75 0.72 0.84 0.68 0.55 0.77 0.97 0.75
Qwen2.5-72B-Instruct 0.77 0.77 0.72 0.85 0.69 0.58 0.78 0.97 0.77
Llama-3.1-8B-Instruct 0.72 0.73 0.72 0.82 0.66 0.55 0.75 0.93 0.74
Qwen2.5-7B-Instruct 0.74 0.74 0.74 0.83 0.64 0.55 0.76 0.95 0.74
Average 0.75 0.75 0.72 0.84 0.67 0.57 0.77 0.96 0.75
MLLMs Single GPT-4o 0.78 0.82 0.76 0.90 0.64 0.53 0.24 0.60 0.66
Llama-3.2-90B-V-Instruct 0.77 0.68 0.64 0.76 0.51 0.34 0.10 0.01 0.48
Qwen2-VL-72B-Instruct 0.77 0.78 0.72 0.88 0.43 0.32 0.15 0.16 0.53
Llama-3.2-11B-V-Instruct 0.77 0.64 0.60 0.68 0.26 0.21 0.02 0.02 0.40
Qwen2-VL-7B-Instruct 0.66 0.64 0.71 0.83 0.39 0.30 0.23 0.15 0.49
Average 0.75 0.71 0.69 0.81 0.53 0.43 0.22 0.19 0.54
Multi GPT-4o 0.77 0.78 0.72 0.86 0.65 0.56 0.81 0.97 0.76
Llama-3.2-90B-V-Instruct 0.72 0.70 0.67 0.74 0.55 0.43 0.62 0.94 0.67
Qwen2-VL-72B-Instruct 0.75 0.74 0.71 0.80 0.63 0.52 0.73 0.94 0.73
Llama-3.2-11B-V-Instruct 0.72 0.69 0.70 0.69 0.47 0.34 0.34 0.81 0.59
Qwen2-VL-7B-Instruct 0.70 0.75 0.73 0.77 0.54 0.42 0.51 0.94 0.67
Average 0.73 0.73 0.70 0.77 0.57 0.46 0.62 0.92 0.69
Table 2: The overall experiment results on M2RAG task. Flu., Rel., CP. and Faith. represent fluency, relevance, context precision and faithfulness, respectively. Coher., Help., Ref., and Recall represent image coherence, helpfulness, reference and recall, respectively. The highest scores for each group are highlighted in bold, and the highest average scores for 4 groups are highlighted in red.
Using the proposed baseline methods and evaluation metrics, we conduct comprehensive experiments on the benchmark dataset with a range of representative models, including GPT-4o, the Llama-3.1 series, Qwen2.5 series, Llama-3.2-Vision series, and Qwen2-VL series [^6] Please refer to Appendix E for more details.. In this section, we analyze the experimental results presented in Table 2 from the following perspectives.
@@
As shown in Table 2, it can be found that the most capable open-source and closed-source foundation models for the M2RAG task are Qwen2.5-72B-Instruct and GPT-4o (0.77), respectively, under the experimental setup of multi-stage approach with LLM backbones. Although the performance gap between these two models are negligible with single-stage approach, there is still a noticeable gap when it comes to single-stage approach(0.68 > 0.66). For MLLM backbones, GPT-4o significantly outperforms all closed-source models. For example, Qwen2-VL-72B-Instruct is much worse than GPT-4o by using the single-stage approach (0.53 < 0.66). Besides, there is still significant room for improvement in the performance of the most capable GPT-4o model (0.77 < 1.0) in the M2RAG task.
@@
Significant scaling phenomena are observed in most experimental setups, with model performance improving as model size increases. However, a notable exception is found in the single-stage approach with LLMs. For instance, Llama-3.1-8B-Instruct slightly outperforms Llama-3.1-70B-Instruct (0.60 > 0.57) under this approach. Rigorous human annotation reveals that this issue arises because larger models tend to select fewer images in the single-stage approach. It should be noted that many of the excluded images are beneficial for the response, and their omission results in performance degradation. For example, the Image Recall scores of Llama-3.1-70B-Instruct and Qwen-2.5-72B-Instruct are significantly lower than those of Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct (0.22 < 0.62 and 0.52 < 0.63, respectively). This phenomenon highlights a limitation in the image selection behavior of larger models, which warrants further investigation in future work.
Joint Modeling versus Separate Modeling
It can be observed that the separate modeling method is much worse than all of our proposed approaches. Specifically, our worst approach, i.e., single-stage with MLLMs (GPT-4o) still outperforms it (0.66 > 0.63). This phenomenon demonstrates that jointly modeling the relationship and interactive information between multi-modal data is crucial for the M2RAG task.
Comparison among Approaches
As shown in the “Average” rows in Table 2, it can be found that the multi-stage approach significantly outperforms the single-stage approaches for LLMs models (0.75 > 0.62) and MLLMs models (0.69 > 0.54). Besides, multi-stage approach introduce more relevant images, enhancing the information density and readability of the generated answers. Specifically, the Image Recall and Helpfulness fine-grained scores of multi-stage approaches are higher than the single-stage approach.
Moreover, the LLMs significantly outperforms the MLLMs in M2RAG tasks, no matter the single-stage or multi-stage is used. For example, the average overall score of LLMs are better than MLLMs (0.71 > 0.64). This observation demonstrates that the capabilities of existing MLLMs are not yet capable enough to accomplish the complex M2RAG tasks.
Model Approach Prompt Completion Total
LLMs Single 7797 767 8563
Multi 16947 2252 19199
MLLMs Single 11454 553 12007
Multi 24078 1707 25785
Table 3: Average token usage for different approaches.
Furthermore, we also compute the average cost, i.e., the average prompt and completion tokens of GPT-4o, in Table 3, and we could summarize two conclusions: (1) the cost multi-stage approach is over 2 times that of single-stage; (2) the cost of MLLMs is approximately 1.3-1.4 times that of LLMs. Therefore, considering that LLMs generally performs better than MLLMs in M2RAG, using single-stage with LLMs as backbone model is currently the most cost-effective approach.

### v2

@@ 4.1.3 Element Evaluation
As described in Section 3, we need to perform In-Doc retrieval to select top-$k$ relevant text and image elements for efficient generation. To achieve this goal, we utilize DeepSeek-Chat DeepSeek-AI (2024) and MiniCPM-V-2.6 Yao et al. (2024) for evaluation, as shown in Figure 2 (Step 3). Our preliminary manual annotations reveal that these methods demonstrate strong correlations with human evaluations in measuring correlation and significantly outperforms the traditional embedding-based models like CLIP Radford et al. (2021), thereby resulting in better quality of our benchmark [^4] Please refer to Appendix C for more details..
In summary, given these data resources, we sample 75 queries for each category (other than Others), resulting in 750 queries for the benchmark dataset.
4.2 Evaluation Metrics
To examine the performance of LLMs and MLLMs on our proposed benchmark, we introduce four text-modal and four multi-modal metrics. Since the generated responses are open-ended, most evaluation metrics are implemented by prompting advanced GPT-4o model in a reference-free manner Lan et al. (2024b).
Text-modal Metrics
Following prior works Zhu et al. (2024), we evaluate the quality of textual generations using both linguistic and RAG metrics Es et al. (2023); Lan et al. (2024a): (1) Fluency assesses the linguistic quality of model-generated text, ensuring the grammatical correctness, coherence, and readability Su et al. (2022); Es et al. (2023); (2) Relevance evaluates how well the model-generated textual content aligns with the given user query Lan et al. (2022); Es et al. (2023); (3) Context Precision measures the proportion of relevant chunks in $K_{\text{In-Doc}}$ by analyzing the overlap of their key concepts Es et al. (2023). (4) Faithfulness measures the accuracy of generations in representing information from $K_{\text{In-Doc}}$, focusing on factual alignment to avoid fabricated or misleading details Es et al. (2023); Sun et al. (2024).
Multi-modal Metrics
Unlike evaluation metrics that focus solely on image quality Tu et al. (2024), our work evaluates the interplay between text and images, covering four key aspects: (1) Image Coherence examines the logical and coherent alignment of images with their surrounding text, ensuring that the visual content enhances and complements the textual narrative Zhu et al. (2024); (2) Image Helpfulness evaluates the contributions of images to the user’s understanding of the text, assessing whether visuals provide additional insights, clarify complex ideas, or support textual details Zhu et al. (2024); (3) Image Reference verifies the appropriateness and alignment between images and their textual references. Inadequate or incorrect references of images lead to the confusion during reading; (4) Image Recall measures the proportion of highly relevant, informative, and important images incorporated into the generations. These crucial images are evaluated and collected by Element Evaluation (Section 4.1.3).
Overall Score Computation
We compute the overall score by taking the average of all text-modal and multi-modal metrics, reflecting the overall capabilities of models. Scores are scaled to [0, 100], with higher values indicating superior performance. For additional details on the specific prompts used for each metric, please refer to Appendix E.
4.3 Generation Strategy for M2RAG
Previous work MuRAR Zhu et al. (2024) generates and integrates text and images independently. We refer to this approach as the separate strategy, which neglects interactions between multi-modal data during text generation. To address this limitation, we propose two joint modeling strategies that explicitly capture the relationships between text and images: the single-stage strategy and the multi-stage strategy, as illustrated in Step 4 of Figure 2.
Single-stage
@@ 4.3 Generation Strategy for M2RAG
requires the model to directly generate a multi-modal output by placing selected images into their corresponding placeholders, using all multi-modal content provided within a single prompt.
Multi-stage
@@
addresses the common limitation of foundation models, which often struggle to process a large number of images simultaneously. It involves three stages: (1) Text Generation: generating a plain text answer based on the multi-modal input; (2) Image Interleaving: dividing the text into segments and prompting LLMs or MLLMs to identify which segments require image insertions for improved readability; (3) Text Refinement: refining each segment by incorporating information from the selected images.
Both LLMs and MLLMs are able to handle M2RAG task with these two strategies. Specifically, LLMs and MLLMs get the prompt $P$ as input to handle M2RAG task, which is constructed using a structured template $T$:
$P=T(G,Q,K_{\text{In-Doc}})$ (2)
where $G,Q,K_{\text{In-Doc}}$ denotes the task guidelines, user query and knowledge base. In the prompt $P$, each image is represented using Markdown format: ![IMAGE_CONTENT](PSEUDO_URL): (1) For MLLMs, IMAGE_CONTENT refers to encoded embeddings of the image, with the entire placeholder positioned where the image originally appears in the context Dubey et al. (2024); Wang et al. (2024a), ensuring coherence between the image and the surrounding text; (2) For LLMs, we convert the image into a detailed textual description as IMAGE_CONTENT, enabling the LLMs to comprehend the semantic information of the image. PSEUDO_URL serves as the identifier for each input image. In the output of the single-stage approach, images are also represented using the aforementioned Markdown format, where PSEUDO_URL indicates the index of the selected image.
4.4 Training Dataset Construction
As another contribution to the research community, we construct a high-quality training dataset to enhance the performance of small-scale but efficient 7B-8B LLMs and MLLMs on the M2RAG task. Specifically, we first utilize the state-of-the-art GPT-4o with multi-stage strategy to construct 3K triplets of ($Q$, $K_{\text{In-Doc}}$, $r$). [^5] Experimental results in Section 5.2 suggest this setup leads to the best performance. Then, our proposed multi-modal metrics are used to filter the high-quality samples. We do not utilize the text-modal metrics, since the time cost of Context Precision and Faithfulness metrics on 3K long generations are huge [^6] Please refer to Section Limitations for more explanation., and we plan to improve the efficiency of text-modal evaluation metrics and update the training data in future work. Finally, the dataset consists of 1.6K instances. More implementation details and statistical information of our training dataset could be found in Appendix D.
@@ 5 Experimental Results
We conduct the comprehensive evaluation results to demonstrate: (1) Reliability of our designed multi-modal automatic evaluation metrics (Section 5.1); (2) A landscape of advanced LLMs and MLLMs performance (Section 5.2); and (3) Effectiveness of our training dataset (Section 5.3).
5.1 Validate Reliability of Evaluation Metrics
Metric Human v.s. Max Min Avg.
Image Coherence MLLM 0.56 0.36 0.48
Human 0.55 0.46 0.51
Image Helpfulness MLLM 0.67 0.48 0.57
Human 0.61 0.47 0.53
Image Reference MLLM 0.61 0.43 0.53
Human 0.67 0.54 0.59
Table 1: The Spearman’s correlation between MLLMs and human annotators. Max, Min and Avg. represents the maximum, minimum and average correlations with all human annotators except for themselves.
Since the reliability of text-modal metrics have been proven in RAG scenarios Es et al. (2023), we mainly focus on validating the reliability of our proposed MLLM-based multi-modal metrics: (1) Image Coherence; (2) Image Helpfulness; and (3) Image Reference. In practice, we randomly select 200 samples from the benchmark labeled by each metric, and three annotators independently scored these samples using the same scoring criteria as the model. Specifically, we calculate the Spearman correlation coefficients between two scorers, each can be human annotators or MLLMs. Table 1 reveals that MLLMs achieve comparable performance to inner-correlation of human annotators. GPT-4o is even better than human average on helpfulness metric (0.57 > 0.53). These results indicate that our proposed multi-modal evaluation metrics are good proxy for human annotators.
5.2 Overall Experimental Results
Model Type Generation Strategy Model Text-modal Metrics Multi-modal Metrics Overall
@@ 5.2 Overall Experimental Results
Reasoner Single OpenAI o3-mini 80.8 89.5 91.9 90.0 73.3 63.2 64.6 79.1 79.1
Multi OpenAI o3-mini 79.6 80.2 94.4 83.8 69.7 60.6 78.0 100.0 80.8
LLMs Separate GPT-4o Zhu et al. (2024) 84.5 88.6 95.0 89.6 58.3 50.4 43.8 82.8 74.1
Single GPT-4o 82.9 88.3 95.7 88.8 72.2 63.8 32.2 75.8 74.9
DeepSeek-V3 81.2 87.9 91.9 87.4 63.6 52.7 36.8 87.3 73.6
Llama-3.1-70B-Instruct 81.2 86.5 93.6 83.8 62.7 55.9 23.8 66.0 69.2
Qwen2.5-72B-Instruct 82.5 89.2 95.8 87.2 66.3 59.6 36.6 78.6 74.5
Llama-3.1-8B-Instruct 79.8 85.2 91.0 81.1 54.1 46.6 26.3 79.5 68.0
Qwen2.5-7B-Instruct 74.2 79.4 88.9 82.1 55.9 48.4 22.2 80.0 66.4
Average 80.3 86.1 92.8 85.1 62.5 54.5 29.6 77.8 71.1
Multi GPT-4o 82.5 87.5 96.6 81.8 73.3 65.8 81.5 100.0 83.6
DeepSeek-V3 80.5 83.0 95.0 85.9 70.0 63.1 79.1 100.0 82.1
Llama-3.1-70B-Instruct 80.7 87.6 97.3 83.9 73.0 63.6 76.8 99.2 82.8
Qwen2.5-72B-Instruct 83.8 88.6 98.4 83.1 73.4 64.5 78.5 99.9 83.8
Llama-3.1-8B-Instruct 81.3 86.5 95.8 82.1 72.1 62.0 75.1 97.9 81.6
Qwen2.5-7B-Instruct 82.8 87.0 95.3 78.5 71.3 62.4 77.9 99.0 81.8
Average 81.9 86.7 96.4 82.5 72.2 63.6 78.1 99.3 82.6
MLLMs Single GPT-4o 81.4 88.6 95.2 88.0 69.8 61.0 25.1 81.5 73.8
Step-1o 84.1 87.4 94.9 90.7 65.9 53.8 44.6 84.7 75.8
Llama-3.2-90B-V-Instruct 82.5 73.9 77.4 79.0 55.7 51.4 18.6 58.0 62.1
Qwen2-VL-72B-Instruct 82.5 85.3 88.3 87.9 50.9 42.1 18.8 62.6 64.8
Llama-3.2-11B-V-Instruct 83.0 71.4 73.0 70.7 37.8 20.6 1.1 58.0 51.9
Qwen2-VL-7B-Instruct 69.0 70.6 79.0 72.9 44.7 39.3 23.6 61.5 57.6
Average 80.4 79.5 84.6 81.5 54.1 44.7 22.0 67.7 64.3
Multi GPT-4o 81.6 88.2 96.6 81.0 72.2 63.5 80.6 100.0 83.0
Step-1o 80.5 82.5 96.7 89.8 71.6 63.7 81.6 97.6 83.0
Llama-3.2-90B-V-Instruct 80.9 82.8 92.0 75.5 60.5 51.0 62.7 98.1 75.4
Qwen2-VL-72B-Instruct 83.0 86.5 94.3 79.0 68.0 58.8 71.2 98.8 79.9
Llama-3.2-11B-V-Instruct 82.8 80.7 87.8 69.9 52.7 42.7 34.4 86.2 67.1
Qwen2-VL-7B-Instruct 80.6 86.7 93.1 79.2 60.3 49.1 49.9 96.0 74.4
Average 81.6 84.5 93.4 79.1 64.2 54.8 63.4 96.1 77.1
Fine-tuned Models Single Llama-3.1-8B-Instruct 74.9 76.5 96.1 82.4 66.3 60.6 75.2 97.0 78.6
Qwen2.5-7B-Instruct 75.3 77.6 94.9 80.6 66.0 58.8 73.9 87.6 76.9
Qwen2-VL-7B-Instruct 57.2 74.3 91.4 71.4 65.4 57.4 73.5 95.0 73.2
Table 2: The overall experiment results on M2RAG task. Flu., Rel., CP. and Faith. represent fluency, relevance, context precision and faithfulness, respectively. Coher., Help., Ref., and Recall represent image coherence, helpfulness, reference and recall, respectively. The highest scores for each group are highlighted in bold, and the highest scores for all settings are highlighted in red.
We conduct comprehensive experiments on our benchmark with 12 representative models, including OpenAI o3-mini, GPT-4o, DeepSeek-V3, Step-1o, Llama-3.1 & Llama-3.2-Vision series, Qwen2.5 & Qwen2-VL series (see Appendix F for details). Due to the high evaluation cost, we sample 200 queries from the benchmark dataset by category to compare the performance of all models. The results are shown in Table 2 , and we can observe several findings.
@@
Table 2 indicates that the most capable open-source and closed-source LLMs for the M2RAG task are Qwen2.5-72B-Instruct and GPT-4o, using a multi-stage strategy with LLM backbones. Among MLLM backbones, GPT-4o and Step-1o outperform all open-source models. For example, Step-1o surpassing Qwen2-VL-72B-Instruct by a large margin (83.0 > 79.9). The reasoner OpenAI o3-mini significantly outperforms other models in the single-stage approach. Despite these strong performances, there remains significant room for improvement, as even the best-performing model falls short of the ideal score (83.8 < 100.0).
@@
A clear scaling trend is observed, where larger open-source models generally yield better results, except in the single-stage approach with LLMs, where Llama-3.1-8B-Instruct surprisingly outperforms Llama-3.1-70B-Instruct in Image Recall (79.5 > 66.0). This discrepancy arises because larger models tend to select fewer images, some of which are beneficial, leading to degraded performance.
Comparison between Separate and Joint Modeling
Regarding modeling approaches, separate modeling Zhu et al. (2024) proves significantly weaker than joint modeling, with even the worst-performing joint approach (GPT-4o, 74.9) surpassing separate modeling (74.1), highlighting the importance of modeling multi-modal interactions.
Comparison among Strategies
Multi-stage approaches consistently outperform single-stage methods for both LLMs (82.6 > 71.1) and MLLMs (77.1 > 64.3), as they introduce more relevant images, improving information density and readability. Additionally, open-source LLMs significantly outperform MLLMs across both modeling paradigms, with an overall score advantage (76.8 > 70.7), indicating that current open-source MLLMs still struggle with the complexity of M2RAG tasks.
Notably, while being constrained by evaluation costs, to investigate whether these findings remain valid as the test data scale increases, we selected several representative models and evaluated them on all 750 queries. The results, presented in Appendix I, confirm that the findings remain valid.
5.3 Improvements from Fine-tuning
We further conduct the supervised fine-tuning to improve the 7B-8B LLMs and MLLMs capabilities in M2RAG task. Please refer to Appendix H for more fine-tuning implementation details. The evaluation results are illustrated in Table 2. It can be observed that our fine-tuned LLMs and MLLMs exhibit significant improvements on all multi-modal evaluation metrics and the overall score. For example, the average improvements on the overall score is 19.2%. Since the text-modal metrics are not utilized for constructing training dataset, our fine-tuned models underperforms the baselines on some textual evaluation metrics, particularly Fluency.